# PhysE-Inv: A Physics-Encoded Inverse Modeling approach for Arctic Snow Depth Prediction

### v1

@@ Abstract
The accurate estimation of Arctic snow depth ($h_{s}$) remains a critical time-varying inverse problem due to the extreme scarcity and noise inherent in associated sea ice parameters. Existing process-based and data-driven models are either highly sensitive to sparse data or lack the physical interpretability required for climate-critical applications. To address this gap, we introduce PhysE-Inv, a novel framework that integrates a sophisticated sequential architecture, an LSTM Encoder-Decoder with Multi-head Attention and physics-guided contrastive learning, with physics-guided inference.Our core innovation lies in a surjective, physics-constrained inversion methodology. This methodology first leverages the hydrostatic balance forward model as a target-formulation proxy, enabling effective learning in the absence of direct $h_{s}$ ground truth; second, it uses reconstruction physics regularization over a latent space to dynamically discover hidden physical parameters from noisy, incomplete time-series input. Evaluated against state-of-the-art baselines, PhysE-Inv significantly improves prediction performance, reducing error by 20% while demonstrating superior physical consistency and resilience to data sparsity compared to empirical methods. This approach pioneers a path for noise-tolerant, interpretable inverse modeling, with wide applicability in geospatial and cryospheric domains.
@@ 1 Introduction
Snow depth ($h_{s}$) exerts a first-order control on Arctic sea ice thickness, yet reliable observations of it remain remarkably scarce. For instance, widely used reanalysis products such as ERA5 do not provide direct measurements of snow depth over sea ice (Hersbach et al., 2020), despite its well-established influence on sea ice thermodynamics (Li et al., 2024). Compounding this data gap, the most available driving observations, such as snow density ($\rho_{s}$), are inherently noisy and sparse due to measurement complexities. To address this gap, we exploit Physics-encoded Neural Networks (PeNNs) (Faroughi et al., 2022), a class of models widely applied in scientific analysis to implement inverse modeling. However, a standard PeNN approach is insufficient to overcome the core issue of extreme data scarcity and input noise. Therefore, our work introduces a novel framework that not only encodes sea ice physics derived from the hydrostatic balance equation (Kwasniok, 2022) but, critically, leverages this equation as a target-formulation proxy to enable the estimation of hidden parameters and yield physically consistent prediction of snow depth ($h_{s}$) from the noisy observational data.
@@ 3.1 Dataset
The dataset comprises ERA5 reanalysis data Hersbach et al. (2020) from the European Centre for Medium-Range Weather Forecasts (ECMWF). ERA5 ingests a wide variety of observational data, a significant portion of which comes from remote sensing instruments Hersbach et al. (2020). We acquired spatiotemporal data from January 1, 1995, to 2024 (10,958 time steps), specifically for the central Arctic Ocean region enclosed by the highlighted orange boundary in Figure 4, which roughly corresponds to the latitude range of approximately 70N to 85N. The data has a spatial resolution of 0.25 $\times$ 0.25 (approximately 25 km). It includes key parameters related to snow depth and sea ice thickness: snow albedo, snow density, and sea ice concentration. We therefore process these variables using a proxy model derived from the hydrostatic balance equation to generate target proxy data that is eventually used in the inversion process to estimate hidden characteristics. The implementation is available here.

### v2

@@ Abstract
The accurate estimation of Arctic snow depth remains a critical time-varying inverse problem due to the extreme scarcity and noise inherent in associated sea ice parameters. Existing process-based and data-driven models are either highly sensitive to sparse data or lack the physical interpretability required for climate-critical applications. To address this gap, we introduce PhysE-Inv, a novel framework that integrates a sophisticated sequential architecture, an LSTM Encoder-Decoder with Multi-head Attention and physics-guided contrastive learning, with physics-guided inference.Our core innovation lies in a surjective, physics-constrained inversion methodology. This methodology first leverages the hydrostatic balance forward model as a target-formulation proxy, enabling effective learning in the absence of direct $h_{s}$ ground truth; second, it uses reconstruction physics regularization over a latent space to dynamically discover hidden physical parameters from noisy, incomplete time-series input. Evaluated against state-of-the-art baselines, PhysE-Inv significantly improves prediction performance, reducing error by 20% while demonstrating superior physical consistency and resilience to data sparsity compared to empirical methods. This approach pioneers a path for noise-tolerant, interpretable inverse modeling, with wide applicability in geospatial and cryospheric domains.
@@ 1 Introduction
Snow depth exerts a first-order control on Arctic sea ice thickness, yet reliable observations of it remain remarkably scarce. For instance, widely used reanalysis products such as ERA5 do not provide direct measurements of snow depth over sea ice (Hersbach et al., 2020), despite its well-established influence on sea ice thermodynamics (Li et al., 2024). Compounding this data gap, the most available driving observations, such as snow density ($\rho_{s}$), are inherently noisy and sparse due to measurement complexities. To address this gap, we exploit Physics-encoded Neural Networks (PeNNs) (Faroughi et al., 2022), a class of models widely applied in scientific analysis to implement inverse modeling. However, a standard PeNN approach is insufficient to overcome the core issue of extreme data scarcity and input noise. Therefore, our work introduces a novel framework that not only encodes sea ice physics derived from the hydrostatic balance equation (Kwasniok, 2022) but, critically, leverages this equation as a target-formulation proxy to enable the estimation of hidden parameters and yield physically consistent prediction of snow depth ($h_{s}$) from the noisy observational data.
@@ 3.1 Dataset
This study utilizes the ERA5 atmospheric reanalysis dataset from the European Centre for Medium-Range Weather Forecasts (ECMWF) Hersbach et al. (2020). ERA5 integrates diverse observational data, primarily from satellite-based remote sensing instruments, to provide consistent daily climate analysis. We extracted spatiotemporal data for the central Arctic Ocean (approximately 70N to 85N) covering the period from January 1, 2019, to December 31, 2023, totaling 1,826 time steps.The data maintains a high spatial resolution of 0.25 $\times$ 0.25 ($\approx$25 km), resulting in a standard grid size of 448 $\times$ 304 pixels. To prepare the dataset for the inversion process, we applied spatial averaging and normalization to key parameters, including snow albedo, snow density, and sea ice concentration. These variables were subsequently processed through a proxy model based on the hydrostatic balance equation to estimate the target hidden characteristics of the sea ice. The implementation is available here.